OpenStudy (anonymous):

Solve: -2x - 15 = 6x + 9

OpenStudy (ipwnbunnies):

Get the x terms on one side and the constant numbers on the other side. Solve for x.

OpenStudy (ipwnbunnies):

-15 - 9 = 6x + 2x --> Simplify and solve for x.

OpenStudy (anonymous):

\[-2x-15=6x+9\]

OpenStudy (anonymous):

\[-2x-6x=9+15\]

OpenStudy (anonymous):

\[-8x=24\]

OpenStudy (anonymous):

\[x=24/-8\]

OpenStudy (anonymous):

\[x=3\]

OpenStudy (anonymous):

i hope u understood it